A researcher wishes to​ estimate, with

99%

​confidence, the population proportion of adults who say chocolate is their favorite ice cream flavor. Her estimate must be accurate within

two​%

of the population proportion.

​(a) No preliminary estimate is available. Find the minimum sample size needed.

​(b) Find the minimum sample size​ needed, using a prior study that found that

44%

of the respondents said their favorite flavor of ice cream is chocolate.

​(c) Compare the results from parts​ (a) and​ (b).

​(a) What is the minimum sample size needed assuming that no prior information is​ available? n=?

(round up to nearest whole number if needed)

​(​b)

What is the minimum sample size needed using a prior study that found that

44%

of the respondents said their favorite ice cream flavor is​ chocolate? n=

(round p to nearest whole number if needed)

(c) How do the results from​ (a) and​ (b) compare?

A. having an estimate of the population proportion raises the minimum sample size needed.

B. having an estimate of the population proportion has no effect on the minimum sample size needed.

C. having an estimate of the population proportion reduces the minimum sample size needed.
Answer #1

Sample size = Z2 * p(1 -p) / E2

a)

When prior estimate for proportion is not available then p = 0.50

Sample size = 2.57582 * 0.50 * 0.50 / 0.022

= 4146.72

n = 4147 (Rounded up to nearest integer)

b)

Given, p = 0.44

Sample size = 2.57582 * 0.44 * 0.56 / 0.022

= 4087.003

Sample size = 4088 (Rounded up to nearest integer)

c)

Having an estimate of the population proportion reduces the minimum sample size needed.